The amount of Jens monthly phone bill is normally distributed with a mean of $50 and a standard deviation of $9. What percentage of her phone bills are between $23 and $77?
Answer #1

Eolution Given deviation standard Mean l=50, as o=9 Z C *P( 28-50 P (234 XL 77) = p (2354 < am foc 77-4) <**,50 =pl -3 <2<3)